the product of -3 and 4 plus a number is -33
Answer by ewatrrr(24785)

 
Hi,
the product of -3 and 4 plus a number is -33
-3(n+4) = -33
n+4 = 11
n = 7
CHECKING our answer**
-3(7 + 4) = -3(11) = -33
Wish You the Best in your Studies.